# M. Hashem Pesaran

### v1

@@ 5.1 Introduction
We apply the procedures proposed above to the problem of forecasting quarterly UK inflation at horizons $h=1,2$ and $4$. Given this choice of target variable, we need to choose the samples used for variable selection and estimation, the pre-selected variables to include in $\boldsymbol{z}_{t},$ the variables to include in the active set $\mathcal{S}_{K,t}=\left\{x_{jt},j=1,2,...,K\right\}$, and the variable selection techniques to use.
@@
The target variable is average annual UK inflation, labelled DPUK4, defined as $\pi_{t+h}=100\times\log(p_{t+h}/p_{t+h-4}),$ where $p_{t}$ is the UK consumer price index taken from the IMF International Financial Statistics. Two sets of variables are considered for inclusion in $\boldsymbol{z}_{t}$. The first set, which we label AR2, includes lags of the target variable $\pi_{t},\pi_{t-1}$ (or equivalently $\pi_{t}$ and $\Delta\pi_{t}).$ Given the importance we attach to global variables and network effects, the second set, which we label ARX2, also includes $\pi_{t}^{\ast}$, $\pi_{t-1}^{\ast}$ (or equivalently $\pi_{t}^{\ast}$ and $\Delta\pi_{t}^{\ast})$ where $\pi_{t}^{\ast}$ is a measure of UK specific foreign inflation constructed using UK trade weights with the other countries. [^8] Specifically, $\pi_{t}^{\ast}=\sum_{j}w_{j}\pi_{jt},$ where $\pi_{jt}$ is the inflation rate in country $j$ and $w_{j}$ is the trade weight of country $j$ with UK.
@@ 5.2 Active set
We now turn to the choice of the members of the active set, $\mathcal{S}_{K,t}=\left\{x_{jt},j=1,2,...,K\right\},$ some of which may be included in $\boldsymbol{z}_{t}.$ While our focus is on forecasting not on building a coherent economic model, our choice of the covariates in the active set is motivated by the large Phillips curve literature, which suggests important roles for demand, supply and expectations variables. The demand and supply variables in the active set are both domestic and foreign from both product and labour markets. Expectations are captured by financial variables. Interaction terms were not included, but the non-linearities that have been investigated in the literature may be picked up by the latent foreign variables. These are represented by $y_{it}^{\ast}$ terms, UK specific measures of foreign inflation and output. Explicit measures of $\mathbf{f}_{it}$ and $\mathbf{g}_{t}$ did not seem to be needed.
@@ 5.3.1 Variable selection procedures
In the OCMT critical value function, $c_{p}(K,\delta)=\Phi^{-1}\left(1-\frac{p}{2K^{\delta}}\right),$ we set $p=0.05$ and $\delta=1$. With $K=52$ this means that we only retain variables with t-ratios (in absolute value) exceeding $c_{0.05}(52,1)=3.3$. We also experimented with setting $\delta=1.5,$ which yields, $c_{0.05}(52,1.5)=3.82$. The results were reasonably robust and we focus on the baseline choice of $\delta=1,$ also recommended by CKP.

### v2

@@ 5.1 Introduction
We apply the procedures proposed above to the problem of forecasting quarterly UK inflation at horizons $h=1,2$ and $4$. The target variable is the headline rate, average annual UK inflation, which is also forecast by the Bank of England. It is labelled DPUK4, defined as $\pi_{t+h}=100\times\log(p_{t+h}/p_{t+h-4}),$ where $p_{t}$ is the UK consumer price index taken from the IMF International Financial Statistics. Forecasting annual rates of inflation at quarterly frequencies is subject to the overlapping observations problem when $h>1$, and it is important that the pre-selected variables in $\boldsymbol{z}_{t}$, or the variables include in the active set $\mathcal{S}_{K,t}=\left\{x_{jt},j=1,2,...,K\right\}$ are all pre-determined (known) at time $t$. Furthermore, as discussed earlier, the variables selected for forecasting inflation at different horizons need not be the same, and the selected variables are also likely to change over time.
@@
Two sets of variables are considered for inclusion in $\boldsymbol{z}_{t}$. The first set, which we label AR2, includes lags of the target variable $\pi_{t},\pi_{t-1}$ (or equivalently $\pi_{t}$ and $\Delta\pi_{t}).$ Given the importance we attach to global variables and network effects, the second set, which we label ARX2, also includes $\pi_{t}^{\ast}$, $\pi_{t-1}^{\ast}$ (or equivalently $\pi_{t}^{\ast}$ and $\Delta\pi_{t}^{\ast})$ where $\pi_{t}^{\ast}$ is a measure of UK specific foreign inflation constructed using UK trade weights with the other countries. [^8] Specifically, $\pi_{t}^{\ast}=\sum_{j}w_{j}\pi_{jt},$ where $\pi_{jt}$ is the inflation rate in country $j$ and $w_{j}$ is the trade weight of country $j$ with UK.
@@ 5.2 Active set
We now turn to the choice of the members of the active set, $\mathcal{S}_{K,t}=\left\{x_{jt},j=1,2,...,K\right\},$ some of which may be included in $\boldsymbol{z}_{t}.$ While our focus is on forecasting not on building a coherent economic model, our choice of the covariates in the active set is motivated by the large Phillips curve literature, which suggests important roles for demand, supply and expectations variables. The demand and supply variables in the active set are both domestic and foreign from both product and labour markets. Expectations are captured by financial variables. Interaction terms were not included, but the non-linearities that have been investigated in the literature may be picked up by the latent foreign variables. These are represented by UK specific measures of foreign inflation and output that could be viewed as estimates of $\mathbf{g}_{t}$ that are tailored to the UK in relation to her trading partners.
@@ 5.3.1 Variable selection procedures
In the OCMT critical value function, $c_{p}(K,\delta)=\Phi^{-1}\left(1-\frac{p}{2K^{\delta}}\right),$ we set $p=0.05$ and $\delta=1$. With $K=52$ this means that we only retain variables with t-ratios (in absolute value) exceeding $c_{0.05}(52,1)=3.3$. To allow for possible serieal correlation, we also experimented with setting $\delta=1.5,$ which yields, $c_{0.05}(52,1.5)=3.82$. The results were reasonably robust and we focus on the baseline choice of $\delta=1,$ also recommended by CKP.